OpenStudy (anonymous):

the surface area of a pyramid is 189cm2 the area of the square base is 49cm2. what is the area of one triangular face of the lateral surface? show the steps and explain your work.

OpenStudy (cwrw238):

total area = 49 + 4 * area of each triangular face 4 * area triangles = 189 - 49 area of triangle face = (189 - 49) / 4 cm2